which conic is this
7x^2-12y^2-14x+24y-28=0?

Answer by stanbon(75887)

which conic is this
7x^2-12y^2-14x+24y-28=0?
Rearrange to get:
7(x^2 - 2x + 1) -12(y^2-2y+1) = 28+7-12
7(x+1)^2 - 12(y-1)^2 = 23
Divide thru by 23 to get:
(x+1)^2/(23/7) -(y-1)^2/(23/12) = 1
------------------------
Depending on what you know about conic equations, not all this
work was necessary to see that the equation is that of a hyperbola.
The fact that the x^2 term and the y^2 term have different signs
is enough to show that it is an equation of a hyperbola.
--------------
Cheers,
Stan H.